A man stands on a scale in an elevator that has an upward acceleration a. The scale reads 956 N. When he picks up a 22.7 kg box, the scale reads 1230 N.

The acceleration a.

a =????
Answer #1

initially

W1 = m1 ( g + a)

956 = m1 g + m1 a

finally

w2 = m2 ( g + a)

1230 = (m1 + 22.7) g + m1 a + 22.7 a

1230 = (m1 g + m1 a) + 22.7* 9.8 + 22.7 a

1230 = 956 + 22.7* 9.8 + 22.7 a

a = 2.27 m/s^2

====

Comment before rate in case any doubt, will reply for sure.. goodluck